Circular No. 6999

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

1SAX J0103.2–7209, AX J0051–722, AND XTE J0055–724 G. L. Israel and L. Stella, Osservatorio Astronomico di Roma; S. Cam-

pana and S. Covino, Osservatorio Astronomico di Brera; D. Ricci, SAX Science Data Center; and T. Oosterbroek, Space Science Department, European Space Agency, on behalf of the BeppoSAX Small Magellanic Cloud (SMC) Team, report the discovery of a new x-ray pulsar, designated as 1SAX J0103.2–7209, during a BeppoSAX observation of the SMC on July 27–28: “The source shows nearly sinusoidal pulsations at a period of 345.2 ± 0.1 s. The energy spectrum is well fitted by a power law with photon index 1.0 ± 0.1, corresponding to an absorbed flux of 3 × 10^{−12} erg s^{−1} cm^{−2} (energy band 2–10 keV). The BeppoSAX source position (α = 1^{h}03^{m}13^{s}, δ = −72^{o}09′.3, equinox 2000.0; estimated error radius 40′′) is consistent with the known supernova remnant SNR 0101–724 (Wang and Wu 1992, Ap.J. Suppl. 78, 391) and the Einstein source 1E 0101.5–7225. The latter source was detected in all Einstein, ROSAT, and ASCA pointings since 1979 at a flux level of a few × 10^{−13} erg s^{−1} cm^{−2}, while no pulsations were detected. Multiband optical images obtained with the 1.5-m Danish telescope at the European Southern Observatory (ESO) on 1998 July 24 show that the Be star of m_{V} = 14.8, originally suggested as the optical counterpart of the Einstein source (Hughes and Smith 1994, A.J. 107, 1363), is the only star that shows strong Hα activity in the BeppoSAX error circle. During a BeppoSAX observation of the SMC on July 24–25, we detected x-ray pulsations of 90.65 ± 0.01 and 58.95 ± 0.01 s from the pulsars AX J0051–722 and XTE J0055–724, respectively. The spacing of the Nov. 1997 (IAUC 6777), Mar. 1998 (IAUC 6858), and July 1998 flares observed from AX J0051–722 further suggests an orbital period of ∼ 120 days. The inferred period derivative is ∼ −0.3 ± 0.1 s/yr. The absorbed flux of XTE J0055–724 is ∼ 7 × 10^{−12} erg s^{−1} cm^{−2} (2–10 keV). Optical spectra (1998 July 27; 1.5-m ESO telescope) obtained for the second brightest star (α = 0^{h}54^{m}56^{s}, δ = −72^{o}26′47′′) included in the x-ray error circle of XTE J0055–724 (IAUC 6822) showed strong Hα and Hβ emission lines. Also, a fainter star (α = 0^{h}54^{m}57^{s}, δ = −72^{o}26′48′′) included in the x-ray error circle of XTE J0055–724 showed an Hα emission line. Further optical and x-ray observations of these pulsars are encouraged.”

SUPERNOVA 1998dh IN NGC 7541 Unfiltered CCD magnitude estimates: Aug. 16.96 UT, 14.5 (L. Bos-

chini, Miragolo, Bergamo, Italy); 24.04, 14.9 (S. Moretti, Forlí, Italy).

1998 August 27  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green